OpenStudy (anonymous):

Given the polynomial f(x)= 3x^4-2x+2, find f(3)

OpenStudy (anonymous):

\[f(3) = 3(3)^{4}-2x+2\] =

OpenStudy (anonymous):

3(3)^4 - 2(3) + 2 =

OpenStudy (anonymous):

3*81-6+2

OpenStudy (anonymous):

=239

OpenStudy (anonymous):

To find f(3), means to evaluate the function at x =3....substitute a 3 everywhere there is an x.